find two numbers whose some is 163 such that when the larger number is divided by the smaller number, the quotient is 3 and the remainder is 3.

Answer by lwsshak3(6522)

find two numbers whose some is 163 such that when the larger number is divided by the smaller number, the quotient is 3 and the remainder is 3.
**
let x=smaller number
63-x=larger number
..
(63-x)/x=3+3/x
63-x=3x+3
4x=60
x=15
63-x=48
smaller number=15
larger number=48